# Enhancement of adhesion strength through microvibrations: modeling and experiments

### v1

@@ Abstract
High-frequency micrometrical vibrations have been shown to greatly influence the adhesive performance of soft interfaces, however a detailed comparison between theoretical predictions and experimental results is still missing. Here, the problem of a rigid spherical indenter, hung on a soft spring, that is unloaded from an adhesive viscoelastic vibrating substrate is considered. The experimental tests were performed by unloading a borosilicate glass lens from a soft PDMS substrate excited by high-frequency micrometrical vibrations. We show that as soon as the vibration starts, the contact area increases abruptly and during unloading it decreases following approximately the JKR classical model, but with a much increased work of adhesion. We find that the pull-off force increases with respect to the amplitude of vibration up to a certain saturation level, which appeared to be frequency dependent. Under the hypothesis of short range adhesion, a lumped mechanical model was derived, which, starting from an independent characterization of the rate-dependent interfacial adhesion, predicted qualitatively and quantitatively the experimental results, without the need of any adjustable parameters.
@@
[inst2]organization=Hamburg University of Technology, Department of Mechanical Engineering,addressline=Am Schwarzenberg-Campus 1, city=Hamburg, postcode=21073, country=Germany
@@
The sphere indentation depends on substrate vibration and indenter oscillation
@@ 2.1 Overview of the mechanical model
where $F_{c,\delta=0}=\left(4/3\right)\pi R\Delta\gamma_{0}$. The normal contact force $F_{c}\left(\delta\right)$ is an implicit equation of the indentation $\delta$ through the JKR model Eq. (2,3).
@@ 2.2 Rate-dependent adhesion
Nevertheless, if we limit our interest to soft elastomers and short range adhesion, dissipation may be assumed to be concentrated close to the crack tip (along the contact periphery), and be accounted for by considering an effective work of adhesion $\Delta\gamma_{eff}$. The latter represents an enhancement of the thermodynamic work of adhesion $\Delta\gamma_{0}$ when the crack propagates, and a reduction when the crack heals (Persson and Brener, 2005; Greenwood and Johnson, 1981; Greenwood, 2004; Nazari et al., 2024; Papangelo et al., 2024; Carbone et al., 2022). Several theories have tried to relate the internal dissipation mechanisms of the viscoelastic material to the effective adhesive energy (Persson and Brener, 2005; Greenwood and Johnson, 1981; Carbone et al., 2022). However, the comparison with the experimental data has remained satisfactorily only in a limited range of crack speeds, perhaps because of other nonlinear rate-dependent mechanisms acting within the process zone (Maghami et al., 2024b), only recently considered in theoretical models (Barthel, 2024; Persson, 2024). Nonetheless, there exists a strong experimental evidence that when the crack is propagating, the effective surface energy increases as a function of the crack speed as a power law (Barthel, 2024; Hui et al., 2022), in accordance with the well known phenomenological model firstly introduced by Gent and Schultz (1972) [^1] Due to the time-temperature superposition in viscoelastic behavior of soft materials the reference velocity can be written as $v_{0}=(\kappa a_{T}^{n})^{-1}$ where $\{\kappa,n\}$ are constants with $10<n<$ 1 and $a_{T}$ is the WLF factor that shifts the viscoelastic modulus at different temperatures (Williams et al., 1955).:
@@ 2.3 Dimensionless notation
The system of ODEs in Eq. (16) was solved in MATLAB© using the ode function ode23t, which can handle moderately stiff dynamical problems and it provides the integration results without the use of numerical damping. For a given preload $\widetilde{F}_{c0}$ the initial contact radius $\widetilde{a}_{0}$ and indenter position $\widetilde{x}_{0}$ are found, while it was always assumed that the unloading phase started when the indenter was at rest, hence $\widetilde{x}_{0}^{\prime}=0$.
@@ 3 Model predictions
Before comparing the model predictions against the experimental data, we give here an overview of the system adhesive behavior to varying amplitude and frequency of the imposed microvibrations. In this section the results will be presented in dimensionless form. Otherwise differently stated, it was assumed $\widetilde{k}=0.0774$, $\widetilde{F}_{c0}=-6.547$, $\widetilde{r}=-0.1$ and for the Gent and Schultz parameters $\widetilde{v}_{0}=9.168\times 10^{-4}$ and $n=0.531$.
@@ 3.1 Influence of frequency and amplitude
As we are considering high frequency microvibrations, whose time-scale is much smaller than the time-scale relative to indenter unloading, it is useful to define averaged (over one vibration period $\widetilde{T}=2\pi/\widetilde{f}$) values of the contact force ($\overline{\widetilde{F}}_{c}$) and contact radius ($\overline{\widetilde{a}}$). Under the hypothesis mentioned above, $\overline{\widetilde{F}}_{c}$ coincides with the averaged spring reaction force $\overline{\widetilde{F}}_{k}$:
@@
Figure 2: a) Top panel: contact radius $\widetilde{a}$ versus the spring reaction force $\widetilde{F}_{k}$ in dimensionless form. The black solid line is the JKR curve, the black dashed line is the Hertz contact model, the red solid line is the oscillating $\widetilde{a}$ versus $\widetilde{F}_{k}$ curve, the blue solid line is its average over one vibration period denoted as $\overline{\widetilde{a}}$ vs $\overline{\widetilde{F}}_{k}$. Bottom panel: contact radius $\widetilde{a}$ versus the dimensionless time $\widetilde{t}$ curve (the red line is the oscillating curve, the blue line is its average value over one vibration period). The curves were obtained by fixing $\widetilde{X}_{b}=8$ and $\widetilde{f}=16.62$. b) Contour plot of the pull-off force $\widetilde{F}_{po}$ normalized with respect to the JKR quasi-static value $\widetilde{F}_{JKR}=3/2$ as a function of the dimensionless vibration amplitude $\widetilde{X}_{b}$ and frequency $\widetilde{f}$. c) Time evolution of the sphere indentation $\widetilde{\delta}$ (black lines) and position $\widetilde{x}$ (green lines), substrate vibration $\widetilde{x}_{b}$ (orange lines) and contact radius $\widetilde{a}$ (red lines). The four panels (numbered from 1 to 4) correspond to the locations highlighted in the map in b). The vibration amplitude is $\widetilde{X}_{b}=8$ for every case, while the excitation frequency is: (1) $\widetilde{f}=2.1$, (2) $\widetilde{f}=9.9$, (3) $\widetilde{f}=16.6$, (4) $\widetilde{f}=30.0$.
@@
Having clarified the typical loading-unloading scenario, the effect of $\widetilde{X}_{b}$ and $\widetilde{f}$ on the pull-off force enhancement $\overline{\widetilde{F}}_{po}/\widetilde{F}_{JKR}$ is shown in the contour plot Fig. 2b. In general, within the hypothesis of the indenter not losing contact with the substrate, $\overline{\widetilde{F}}_{po}$ increases when the vibration amplitude is increased, while the behavior with respect to the frequency is influenced by the nonlinear stiffness introduced by the contact, namely $k_{c}$.
Indeed, substituting in the equation of motion (Eq. (5)) the expressions for $\delta$ (Eq. (6)) and its derivatives, in the limit of very high frequency and with low stiffness and damping, which is our case of interest ($\widetilde{k}\ll 1$ and $\zeta\ll 1$), one obtains the resonance frequency of the system at a given mean indentation $\overline{\delta}_{0}$ to be $\omega_{nc}=\sqrt{\left(\left.\frac{dF_{c}\left(\delta\right)}{d\delta}\right|_{\delta=\overline{\delta}_{0}}\right)/m}$ (see Appendix B), where the stiffening effect coming from the contact is at play. For a typical contact radius of $\widetilde{a}=5.1$ one finds $\widetilde{k}_{c}=7.1$ and $\widetilde{f}_{nc}=\omega_{nc}/\omega_{n}\simeq 9.6$ that is consistent with the numerical results found.
Notice that in the JKR model, the relation contact force versus indentation is an implicit equation of the contact radius $\widetilde{a}$, nevertheless an approximation is to consider the stiffness of a flat punch of radius $\widetilde{a}$ that, in our case, would estimate the contact stiffness as $\widetilde{k}_{c}=\left(3/2\right)\widetilde{a}\simeq 7.6$, giving $\widetilde{f}_{nc}\simeq 10$ within the $5\%$ error. Scrutiny of the contour plot in Fig. 2b would also reveal the stiffening behavior of the contact as the locus of the maxima of the pull-off force bends towards the right as $\widetilde{X}_{b}$ increases, as one should expect.
@@
Prior to resonance (case 1), indentation and base displacement move in-phase (consider the convention of signs in Fig. 1) meaning that as the substrate moves the indenter follows. This results in a minimal indentation depth $\widetilde{\delta}$ and in a very small variation of the contact radius $\widetilde{a}$, preventing adhesion enhancement. Case 2 describes the system behavior near the resonance: the indenter vibration this time dominates with respect to the base excitation. This implies $\widetilde{\delta}\approx\widetilde{x}$, which are almost in quadrature of phase with respect to the base excitation $\widetilde{x}_{b}$. Large, high frequency, oscillations of $\widetilde{\delta}$ result in large oscillations of the contact radius, yielding a significant amplification of the adhesive force. By increasing the excitation frequency above the resonance frequency, the indentation and base displacement move out-of-phase (consider the convention of signs in Fig. 1), which signifies that the base vibration is amplified by the oscillator motion resulting in $\widetilde{\delta}>\widetilde{x}_{b}$ (cases 3, 4). Hence, within our modeling assumptions, we have $\lim_{\widetilde{f}\rightarrow+\infty}\widetilde{\delta}(\widetilde{f})=\widetilde{x}_{b}$, resulting in a very mild dependence on the excitation frequency, as it can be clearly seen in Fig. 2b.
Figure 3: Normalized vibration amplitude for the contact radius ($\widetilde{a}_{A}/\overline{\widetilde{a}}$, left y-axis, blue curve) and indentation ($\widetilde{\delta}_{A}/\widetilde{X}_{b}$, right y-axis, orange curve) as a function of the frequency $\widetilde{f}$ for $\widetilde{X}_{b}=8$. The quantities are estimated at the pull-off point.
@@ 3.2 Behaviour at pull-off
To clarify the mechanism behind the adhesion enhancement, Fig. 3 shows how the oscillating parts of the contact radius and indentation just prior to pull-off vary with the excitation frequency, for a fixed amplitude $\widetilde{X}_{b}=8$. The representative quantities $\widetilde{a}_{A}=\left(\max\left(\widetilde{a}\right)-\min\left(\widetilde{a}\right)\right)/2$ and $\widetilde{\delta}_{A}=\left(\max\left(\widetilde{\delta}\right)-\min\left(\widetilde{\delta}\right)\right)/2$, normalized by the mean contact radius $\overline{\widetilde{a}}$ and the substrate variation amplitude $\widetilde{X}_{b}$, respectively, are plotted as function of the frequency. Both curves show the same qualitative behavior, with the oscillating part of the contact radius and indentation increasing at low frequencies, then reaching a peak at the resonance frequency and later decreasing as the frequency is further increased. Notice that the contact area variation is predicted to be quite severe at resonance, being $\widetilde{a}_{A}\approx 0.5\ast\overline{\widetilde{a}}$, while this oscillation steeply drops moving away from the resonance frequency and remaining at high frequency within the $5\%$ of the average contact radius. This is consistent with the experimental observation in Shui et al. (2020). Qualitatively similar conclusions can be drawn for the variation of $\widetilde{\delta}_{A}$. Notice that as the frequency increases, $\widetilde{\delta}_{A}/\widetilde{X}_{b}\rightarrow 1$. At very high frequency $\widetilde{f}\gg\widetilde{f}_{nc}$, a simplified model for the system in Fig. 1 can be considered that neglects the indenter dynamics, which is briefly presented in C.
@@ 4.1 Description of the experimental test-rig
The customized test rig shown in Fig. 4 was used to conduct the vibration-regulated adhesion experiments. The indenter, a borosilicate crown glass plano-convex spherical lens with radius of curvature $R=51.5$ mm, was connected to a uniaxial S-beam load cell (Futek LSB205, natural frequency 1180 Hz, resolution 1 mN). In order to shield the load cell from highly oscillating axial loads during the testing, a compliant rubber band buffer with a stiffness of approximately $320$ N/m was positioned between the load cell and the indenter. Additional weights were placed on top of the indenter to keep the rubber band in tension. A micrometrical motorized linear stage (Physik Instrumente, M-403.2DG) was used to drive the assembly in the direction normal to the PDMS surface.
@@
The dynamic excitation was supplied by two electrodynamic shakers (Dynalabs, DYN-PM-250), that also served as supports for the PMMA transparent beam where the PDMS sample (fixed onto a glass slide) was placed. The sample was positioned right in the center of the beam and two piezoelectric accelerometers (DYTRAN 3055D1, frequency range 1-10000 Hz) where mounted at its sides, allowing to measure the acceleration of the PMMA beam close to the PDMS substrate. The vibration amplitude was then measured by double-integration of the acceleration signal. The measured signals were acquired and elaborated through a data acquisition system (Dewesoft, DAQ system KRYPTON-6XSTG). A high-resolution camera (Alvium 1800 U-2040m, resolution 20 Mpixel, pixel size 2.74 $\mu$m, frame rate 10 fps), captured live images of the contact patch through a right angle optical element positioned beneath the contact. By using a custom code built in Wolfram Mathematica© to post-process the images, the corresponding evolution of the contact radius was determined.
@@
The standard loading protocol was the following: the sample was loaded with a $F_{c0}=-250$ mN preload (loading rate $100$ $\mu$m/s) and a dwell time of $60$ s was allowed for material relaxation. Subsequently, the substrate vibration was switched by starting the electrodynamics shakers and the target amplitude was reached by manually adjusting the amplifier’s gain. An additional $60$ s dwell time was considered in order to stabilize all the acquired signals, e.g. force, vibration amplitude, contact radius. The sample was finally unloaded at $r=-5$ $\mu$m/s. All the experimental tests were performed at room temperature $\approx 20^{\circ}C$.
@@
Figure 5: Characterization of the interface rate-dependent adhesion. The effective work of adhesion $\Delta\gamma_{eff}$ as a function of the crack speed obtained from the experimental data (markers) through Eq. (18). The data points were fitted using Eq. (7) (solid line), resulting in a reference speed of $v_{0}=0.0445$ mm/s and an exponent $n=0.531$. The inset shows a log-log representation of the quantity $(\Delta\gamma_{eff}/\Delta\gamma_{0}-1)$ that highlights the power law scaling.

### v2

@@ Abstract
High-frequency micrometrical vibrations have been shown to greatly influence the adhesive performance of soft interfaces, however a detailed comparison between theoretical predictions and experimental results is still missing. Here, the problem of a rigid spherical indenter, hung on a soft spring, that is unloaded from an adhesive viscoelastic vibrating substrate is considered. The experimental tests were performed by unloading a borosilicate glass lens from a soft PDMS substrate excited by high-frequency micrometrical vibrations. We show that as soon as the vibration starts, the contact area increases abruptly and during unloading it decreases following approximately the JKR classical model, but with a much increased work of adhesion with respect to its thermodynamic value. We find that the pull-off force increases with the amplitude of vibration up to a certain saturation level, which appeared to be frequency dependent. Under the hypothesis of short range adhesion, a lumped mechanical model was derived, which, starting from an independent characterization of the rate-dependent interfacial adhesion, predicted qualitatively and quantitatively the experimental results, without the need of any adjustable parameters.
@@
[…]
@@
The indenter dynamics strongly influences the adhesion strength
@@ 2.1 Overview of the mechanical model
where $F_{c,\delta=0}=\left(4/3\right)\pi R\Delta\gamma_{0}$. The normal contact force $F_{c}\left(\delta\right)$ is an implicit equation of the indentation $\delta$ through the JKR model Eq.s (2,3).
@@ 2.2 Rate-dependent adhesion
Nevertheless, if we limit our interest to soft elastomers and short range adhesion, dissipation may be assumed to be concentrated close to the crack tip (along the contact periphery), and be accounted for by considering an effective work of adhesion $\Delta\gamma_{eff}$. The latter represents an enhancement of the thermodynamic work of adhesion $\Delta\gamma_{0}$ when the crack propagates, and a reduction when the crack heals (Persson and Brener, 2005; Greenwood and Johnson, 1981; Greenwood, 2004; Nazari et al., 2024; Papangelo et al., 2024; Carbone et al., 2022). Several theories have tried to relate the internal dissipation mechanisms of the viscoelastic material to the effective adhesive energy (Persson and Brener, 2005; Greenwood and Johnson, 1981; Carbone et al., 2022). However, the comparison with the experimental data has remained satisfactory only in a limited range of crack speeds, perhaps because of other nonlinear rate-dependent mechanisms acting within the process zone (Maghami et al., 2024b), only recently considered in theoretical models (Barthel, 2024; Persson, 2024). Nonetheless, there exists a strong experimental evidence that when the crack is propagating, the effective surface energy increases as a function of the crack speed as a power law (Barthel, 2024; Hui et al., 2022), in accordance with the well known phenomenological model firstly introduced by Gent and Schultz (1972) [^1] Due to the time-temperature superposition in viscoelastic behavior of soft materials the reference velocity can be written as $v_{0}=(\kappa a_{T}^{n})^{-1}$ where $\{\kappa,n\}$ are constants with $10<n<$ 1 and $a_{T}$ is the WLF factor that shifts the viscoelastic modulus at different temperatures (Williams et al., 1955).:
@@ 2.3 Dimensionless notation
The system of ODEs in Eq. (16) was solved in MATLAB© using the ode function ode23t, which can handle moderately stiff dynamical problems and it provides the integration results without the use of numerical damping. For a given preload $\widetilde{F}_{c0}$ the initial contact radius $\widetilde{a}_{0}$ and indenter position $\widetilde{x}_{0}$ are found, while it was always assumed that the unloading phase started when the indenter was at rest, hence $\widetilde{x}_{0}^{\prime}=0$. The numerical simulations were terminated when, during unloading, the contact radius fell below 0.
@@ 3 Model predictions
Before comparing the model predictions against the experimental data, we give here an overview of the system adhesive behavior to varying amplitude and frequency of the imposed microvibrations. In this section the results will be presented in dimensionless form. Otherwise differently stated, it was assumed $\widetilde{k}=0.0774$, $\widetilde{F}_{c0}=-6.547$, $\widetilde{r}=-0.1$ and for the Gent and Schultz parameters $\widetilde{v}_{0}=9.168\times 10^{-4}$ and $\alpha=0.531$.
@@ 3.1 Influence of frequency and amplitude
As we are considering high frequency microvibrations, whose time-scale is much smaller than the time-scale relative to the indenter unloading, it is useful to define averaged (over one vibration period $\widetilde{T}=2\pi/\widetilde{f}$) values of the contact force ($\overline{\widetilde{F}}_{c}$) and contact radius ($\overline{\widetilde{a}}$). Considering the indenter is slowly unloaded, $\overline{\widetilde{F}}_{c}$ coincides with the averaged spring reaction force $\overline{\widetilde{F}}_{k}$:
Figure 2: a) Top panel: contact radius $\widetilde{a}$ versus the spring reaction force $\widetilde{F}_{k}$ in dimensionless form. The black solid line is the JKR curve, the black dashed line is the Hertz contact model, the red solid line is the oscillating $\widetilde{a}$ versus $\widetilde{F}_{k}$ curve, the blue solid line is its average over one vibration period denoted as $\overline{\widetilde{a}}$ vs $\overline{\widetilde{F}}_{k}$. Bottom panel: contact radius $\widetilde{a}$ versus the dimensionless time $\widetilde{t}$ curve (the red line is the oscillating curve, the blue line is its average value over one vibration period). The curves were obtained by fixing $\widetilde{X}_{b}=8$ and $\widetilde{f}=16.62$. b) Map of the pull-off force $\widetilde{F}_{po}$ normalized with respect to the JKR quasi-static value $\widetilde{F}_{JKR}=3/2$ as a function of the dimensionless vibration amplitude $\widetilde{X}_{b}$ and frequency $\widetilde{f}$. c) Time evolution of the sphere indentation $\widetilde{\delta}$ (black lines) position $\widetilde{x}$ (green lines), substrate vibration $\widetilde{x}_{b}$ (orange lines) and contact radius $\widetilde{a}$ (red lines). The four panels (numbered from 1 to 4) correspond to the locations highlighted in the map in b). The vibration amplitude is $\widetilde{X}_{b}=8$ for every case, while the excitation frequency is: (1) $\widetilde{f}=2.1$, (2) $\widetilde{f}=9.9$, (3) $\widetilde{f}=16.6$, (4) $\widetilde{f}=30.0$.
@@
Having clarified the typical loading-unloading scenario, the effect of $\widetilde{X}_{b}$ and $\widetilde{f}$ on the pull-off force enhancement $\overline{\widetilde{F}}_{po}/\widetilde{F}_{JKR}$ is shown in the map in Fig. 2b. In general the pull-off force $\overline{\widetilde{F}}_{po}$ increases when the vibration amplitude is increased, while the behavior with respect to the frequency is influenced by the nonlinear stiffness introduced by the contact $k_{c}$. At a given indentation $\overline{\delta}_{0}$ the contact stiffness is $k_{c}=\left(dF_{c}/d\delta\right)|_{\delta=\overline{\delta}_{0}}$, which depends on the indentation $\overline{\delta}_{0}$ as the force versus indentation curve is nonlinear. The corresponding dimensionless contact stiffness will be $\widetilde{k}_{c}=k_{c}/(\frac{4}{3}E^{*}a_{r})$. In the numerical simulations the contact model is always satisfied, i.e. there is no contact loss except at the pull-off point where the simulation is truncated. Clearly too large vibration amplitudes may cause detachment of the indenter as soon as the vibrations start. This regime was not investigated here as it may lead to cycles of jump-in and -out of contact including impacts, which we were not interested in.
Indeed, substituting in the equation of motion (Eq. (5)) the expressions for $\delta$ (Eq. (6)) and its derivatives, in the limit of very high frequency and with low apparatus stiffness and damping, which is our case of interest ($\widetilde{k}\ll 1$ and $\zeta\ll 1$), one obtains the resonance frequency of the system at a given mean indentation $\overline{\delta}_{0}$ to be $\omega_{nc}=\sqrt{\left(\left.\frac{dF_{c}\left(\delta\right)}{d\delta}\right|_{\delta=\overline{\delta}_{0}}\right)/m}$ (see Appendix B). For a typical contact radius of $\widetilde{a}=5.1$ one finds $\widetilde{k}_{c}=7.1$ and $\widetilde{f}_{nc}=\omega_{nc}/\omega_{n}\simeq 9.6$ that is consistent with the numerical results found.
Notice that in the JKR model, the relation contact force versus indentation is an implicit equation of the contact radius $\widetilde{a}$, nevertheless an approximation is to consider the stiffness of a flat punch of radius $\widetilde{a}$ that, in our case, would estimate the contact stiffness as $\widetilde{k}_{c}=\left(3/2\right)\widetilde{a}\simeq 7.6$, giving $\widetilde{f}_{nc}\simeq 10$ within the $5\%$ error. Scrutiny of the map in Fig. 2b would also reveal the stiffening behavior of the contact as $k_{c}$ increases with the indentation hence the locus of the maxima of the pull-off force bends towards the right due to the increase of the system resonance frequency $\omega_{nc}$.
@@
Prior to resonance (case 1), indentation and base displacement move in-phase (consider the convention of signs in Fig. 1 meaning that as the substrate moves the indenter follows. This results in a minimal variation of indentation depth $\widetilde{\delta}$ and in a very small variation of the contact radius $\widetilde{a}$, preventing adhesion enhancement. Case 2 describes the system behavior near the resonance: the indenter vibration this time dominates with respect to the base excitation. This implies $\widetilde{\delta}\approx\widetilde{x}$, which are almost in quadrature of phase with respect to the base excitation $\widetilde{x}_{b}$. Large, high frequency, oscillations of $\widetilde{\delta}$ result in large oscillations of the contact radius, yielding a significant amplification of the adhesive force. By increasing the excitation frequency above the resonance frequency ($\widetilde{f}_{nc}\approx 10$), the indentation and base displacement move out-of-phase (consider the convention of signs in Fig. 1), which signifies that the base vibration is amplified by the oscillator motion resulting in $\widetilde{\delta}>\widetilde{x}_{b}$ (cases (3, 4)). Hence, within our modeling assumptions, we have $\lim_{\widetilde{f}\rightarrow+\infty}\widetilde{\delta}(\widetilde{f})=\widetilde{x}_{b}$, resulting in a very mild dependence on the excitation frequency, as it can be clearly seen in Fig. 2b.
Figure 3: Normalized vibration amplitude for the contact radius ($\widetilde{a}_{A}/\overline{\widetilde{a}}$, left y-axis, blue curve) and indentation ($\widetilde{\delta}_{A}/\widetilde{X}_{b}$, right y-axis, orange curve) as a function of the frequency $\widetilde{f}$ for $\widetilde{X}_{b}=8$. The quantities are estimated when the vibrations are switched on and the sphere is subjected to the preload $\widetilde{F}_{c0}=-6.547$.
@@ 3.2 Behaviour at pull-off
To clarify the mechanism behind the adhesion enhancement, Fig. 3 shows how the oscillating parts of the contact radius and indentation (taken as soon as the vibrations are switched on) vary with the excitation frequency, for a fixed amplitude $\widetilde{X}_{b}=8$. The representative quantities $\widetilde{a}_{A}=\left(\max\left(\widetilde{a}\right)-\min\left(\widetilde{a}\right)\right)/2$ and $\widetilde{\delta}_{A}=\left(\max\left(\widetilde{\delta}\right)-\min\left(\widetilde{\delta}\right)\right)/2$, normalized by the mean contact radius $\overline{\widetilde{a}}$ and the substrate vibration amplitude $\widetilde{X}_{b}$, respectively, are plotted as functions of the frequency. Both curves show the same qualitative behavior, with the oscillating part of the contact radius and indentation increasing at low frequencies, then reaching a peak at the resonance frequency and later decreasing as the frequency is further increased. Notice that the contact area variation is predicted to be quite severe at resonance, being $\widetilde{a}_{A}\approx 0.33\ast\overline{\widetilde{a}}$, while this oscillation steeply drops moving away from the resonance frequency and remaining at high frequency within the $1\%$ of the average contact radius. This is consistent with the experimental observation in Shui et al. (2020). Qualitatively similar conclusions can be drawn for the variation of $\widetilde{\delta}_{A}$. Notice that as the frequency increases, $\widetilde{\delta}_{A}/\widetilde{X}_{b}\rightarrow 1$. At very high frequency $\widetilde{f}\gg\widetilde{f}_{nc}$, a simplified model for the system in Fig. 1 can be considered that neglects the indenter dynamics, which is briefly presented in C.
@@ 4.1 Description of the experimental test-rig
The customized test rig shown in Fig. 4 was used to conduct the vibration-regulated adhesion experiments. The indenter, a borosilicate crown glass plano-convex spherical lens with radius of curvature $R=51.5$ mm, was connected to a uniaxial S-beam load cell (Futek LSB205, natural frequency 1180 Hz, resolution 1 mN). In order to shield the load cell from highly oscillating axial loads during the testing, a compliant rubber band buffer was positioned between the load cell and the indenter. Additional weights were placed on top of the indenter to keep the rubber band in tension. The stretched rubber band was $4$ cm long, with a squared cross section of $1\times 1$ mm2. It was connected on both ends (one end to the load cell, the other to the indenter) through steel eye-bolts. The calibration was made by fixing the end connected to the indenter and pulling on the rubber band by moving the upper end connected to the load cell. Force-displacement curves were recorded using the same loading rate of the vibration tests, in a range of forces similar to the one employed in the tests, where the force-displacement curve is approximately linear. A stiffness of $k=320$ N/mm was estimated.
A micrometrical motorized linear stage (Physik Instrumente, M-403.2DG) was used to drive the assembly in the direction normal to the PDMS surface.
@@
The dynamic excitation was supplied by two electrodynamic shakers (Dynalabs, DYN-PM-250), that also served as supports for the PMMA transparent beam where the PDMS sample (fixed onto a glass slide) was placed. The sample was positioned right in the center of the beam and two piezoelectric accelerometers (DYTRAN 3055D1, frequency range 1-10000 Hz) where mounted at its sides, allowing to measure the acceleration of the PMMA beam close to the PDMS substrate. The vibration amplitude was then measured by double-integration of the acceleration signal. The measured signals were acquired and elaborated through a data acquisition system (Dewesoft, DAQ system KRYPTON-6XSTG). A high-resolution camera (Alvium 1800 U-2040m, resolution 20 Mpixel, pixel size 2.74 $\mu$m, frame rate 10 fps, shutter exposure time $5$ ms), captured live images of the contact patch through a right angle optical element positioned beneath the contact. The purpose of the imaging was to measure the averaged values of the contact radius during the whole test. The contact edge was sharp enough to clearly distinguish an average contact circle in most of the tests. Some blur was visible only at large amplitudes or near the pull-off.
By using a custom code built in Wolfram Mathematica© to post-process the images, the corresponding evolution of the contact radius was determined.
@@
The standard loading protocol was the following: the sample was loaded with a $F_{c0}=-250$ mN preload (loading rate $100$ $\mu$m/s) and a dwell time of $60$ s was allowed for material relaxation. Subsequently, the substrate vibration was switched on by starting the electrodynamics shakers and the target amplitude was reached by manually adjusting the amplifier’s gain. An additional $60$ s dwell time was considered in order to stabilize all the acquired signals, e.g. force, vibration amplitude, contact radius. The sample was finally unloaded at $r=-5$ $\mu$m/s. All the experimental tests were performed at room temperature $\approx 20^{\circ}C$.
@@
Figure 5: Characterization of the interface rate-dependent adhesion. The effective work of adhesion $\Delta\gamma_{eff}$ as a function of the crack speed obtained from the experimental data (markers) through Eq. (18). The data points were fitted using Eq. (7) (solid line), resulting in a reference speed of $v_{0}=0.0445$ mm/s and an exponent $\alpha=0.531$. The inset shows a log-log representation of the quantity $(\Delta\gamma_{eff}/\Delta\gamma_{0}-1)$ that highlights the power law scaling.